### v1

@@ 1 Introduction
Stack layouts were introduced by Bernhart and Kainen in 1979 [13] with the motivation to study a “quite natural” edge decomposition technique. A stack layout is a linear layout, that is, the vertices are positioned on a line $\ell$ with the edges being embedded on several half-planes, called pages, delimited by $\ell$. In a stack layout, edges embedded on the same page are forbidden to cross, a rule that contributed to their original designation as book embeddings. An example of stack layout on one page is depicted in Figure 1(b). The minimum number of pages required to obtain a stack layout of a graph is called the stack number. After their introduction, stack layouts attracted notable interest in the field of theoretical computer science due to their capacity to model certain aspects in several domains, including VLSI design [17, 18], fault-tolerant multiprocessing [18, 40], RNA folding [27], and traffic-light control [32]. This led, in particular, to a series of papers investigating upper bounds on the stack number of planar graphs [16, 28, 31] and culminating at an upper bound of four [42]. Bekos et al. [11] and Yannakakis [43] independently proved that this upper bound is tight.
@@
Based on this latter interpretation of stack layouts, Heath, Leighton and Rosenberg [29, 30] proposed to study queue layouts of graphs where the edges occurring on the same page of the linear layout can be represented by a queue. More precisely, edges ending at $v_{i}$ are first dequeued from a queue, after which edges beginning at $v_{i}$ are enqueued during the left-to-right sweep of $\ell$. Thus, edges on the same page are allowed to cross but forbidden to properly nest. An example of a queue layout on one page is shown in Figure 1(c). Surprisingly, these quite non-planar representations still found many practical applications, including scheduling [14], VLSI [34], and, a decade after their inception, 3D crossing-free graph drawing [20, 23]. Similarly to stack layouts, one is interested in minimizing the number of pages required to obtain a queue layout, which is known as the queue number. However, in contrast to the stack number, the queue number of planar graphs remained elusive for several decades during which only sublinear (but non-constant) upper bounds have been found [1, 5, 6, 21]. The first significant difference between planar and non-planar graphs was only demonstrated in 2019 when it was shown that bounded degree planar graphs have bounded queue number [9], whereas non-planar bounded degree graphs were already known to not have this property [41]. Recently, Dujmović et al. [22] proved that the queue number of planar graphs is bounded by a constant. Minor improvements have since been made [10, 25].
@@
We introduce a model called priority queue layout, or simply PQ-layout. This model utilizes priority queues for the storage of edges assigned to the same page, where the edge weights correspond to the priorities (i.e., keys) for the data structure. As for stack and queue layouts, in a PQ-layout all vertices of the graph lie on a horizontal line $\ell$, and the edges are assigned to different priority queues according to a left-to-right sweep of $\ell$. When an edge $e$ is dequeued, it must have the minimum priority among the edges stored in the same priority queue as $e$. See Figure 1(d) for an example, and refer to Section 2 for a formal definition of our model. Note that, unlike stack and queue layouts, our model allows edges in the same page to cross or properly nest, provided that they have suitable weights.
@@ 2 PQ-Layouts
The order of $u$ and $u^{\prime}$ is irrelevant for F.1. For completeness, the resulting three arrangements of the involved vertices are illustrated in Figures 2(a), 2(b) and 2(c). Note that the forbidden pseudo-nesting in Figure 2(b) requires $u=u^{\prime}$. In contrast, the symmetric case where $v=v^{\prime}$ is never forbidden. This is a direct consequence of the asymmetry between the insert and pull operations of priority queues. Thus, unlike for stack and queue layouts, the reverse vertex ordering does not necessarily provide a PQ-layout.
@@
Figure 2: (a)–(c) Arrangements in a PQ-layout that correspond to the Forbidden Configuration F.1. (d) A $k$-inversion, a forbidden arrangement in linear layouts with priority queue number $k-1$.
@@
Some of our proofs utilize configurations where every pair of edges must be assigned to distinct priority queues. Let $\langle G,w\rangle$ be an edge-weighted graph, let $\prec$ be a vertex ordering of $G$, and let $k$ be an integer such that $k\geq 1$. We define a $k$-inversion to be a sequence $(e_{1},\ldots,e_{k})$ of $k$ edges in $G$ with pairwise distinct right endpoints $v_{1},\ldots,v_{k}$, respectively, such that (i) $v_{k}\prec\cdots\prec v_{1}$, where $v_{k}$ and $v_{1}$ are called the left and right end, respectively, (ii) the left endpoint of each of $e_{1},\ldots,e_{k}$ appears to the left of $v_{k}$ in $\prec$, and (iii) $w(e_{1})<\cdots<w(e_{k})$. See Figure 2(d) for an illustration. Any two edges in a $k$-inversion form a forbidden configuration and thus must be assigned to distinct priority queues. In other words, if, for edge weights $w$ of $G$, every vertex ordering of $\langle G,w\rangle$ contains a $k$-inversion, then $\operatorname{pqn}(G,w)\geq k$ and hence $\operatorname{pqn}(G)\geq k$.
@@ 3 Priority Queue Number of Complete Graphs
Theorem 3.1 also provides an upper bound for the (non-separated) priority queue number of $K_{n,n}$. Next, we give a corresponding linear lower bound showing that $\operatorname{pqn}(K_{n,n})\in\Theta(n)$.
@@
For $K_{n,n}$, we define an edge-weight function $w$ that has, at each vertex, for every $i\in\{1,\ldots,n\}$, exactly one incident edge of weight $i$ [^1] In contrast to the proof of Theorem 3.1, we cannot arbitrarily assign weights to edges incident to each vertex $b\in B$ because this may give twice the same edge weight at a vertex of partition $A$. To this end, we partition the edge set into $n$ perfect matchings $M_{1},\ldots,M_{n}$, and, for each edge $e\in M_{i}$, we set $w(e)=i$.
@@
Note that we cannot directly apply Theorem 3.1 to $\Gamma^{\prime}$ because we have already fixed an edge-weight function. Instead, we analyze the possible distribution of the edge weights incident to vertices in $\Gamma^{\prime}$ by a representation as a black-and-white grid; see the appendix for an illustration and a more extensive description. [^3] A similar, although not identical, representation is used by Alam et al. [35] to prove the mixed page number of $K_{n,n}$. Our grid has $\frac{n}{2}$ columns that represent the vertices of the latter partition (in order from left to right as they appear in $\Gamma^{\prime}$) and it has $n$ rows that represent the edge weights $\{1,\dots,n\}$ (in order from bottom to top). A grid cell in column $i$ and row $j$ is colored black if the $i$-th vertex is incident to an edge of weight $j$, and it is colored white otherwise. Observe that, in this grid, a strictly monotonically decreasing path of black cells having length $k$ is equivalent to a $k$-inversion in $\Gamma^{\prime}$. We can show that there always exists such a path of length $\left\lceil\frac{3-\sqrt{5}}{4}n\right\rceil$: the candidates for being the first black cell of the path are those who do not have another black cell in their top left region of the grid. We find these cells, color them white and iteratively repeat this process to find the candidates for being the second, third, etc. black cell of the path. We can prove that we need at least $\left\lceil\frac{3-\sqrt{5}}{4}n\right\rceil$ iterations until the grid is white.
Since $K_{n}$ contains $K_{\lfloor\frac{n}{2}\rfloor,\lfloor\frac{n}{2}\rfloor}$ as a subgraph, Theorem 3.3 immediately implies a linear lower bound on $\operatorname{pqn}(K_{n})$:
@@ 4 Characterizing Graphs with Priority Queue Number 1
We study which graphs have priority queue number 1. We will provide a characterization of the graphs with priority queue number 1 in terms of forbidden minors (Section 4.4). We first show that the family of graphs with priority queue number 1 is minor-closed (Section 4.1). Then, we describe families of graphs with priority queue number 1 (Section 4.2) and a set of forbidden minors (Section 4.3). The characterization of Section 4.4 follows by proving that every graph that does not contain any of the forbidden minors of Section 4.3 falls in one of the families of Section 4.2. This characterization leads to an efficient recognition algorithm.
4.1 PQ-Layouts of Minors
[…]
Let $G=(V,E)$ be a graph with $\operatorname{pqn}(G)=1$ and let $G^{\prime}=G/e_{0}$ be obtained from $G$ by contracting an edge $e_{0}=uv\in E$. Then, $\operatorname{pqn}(G^{\prime})=1$.
Proof 4.2 (Proof Sketch).
We prove $\operatorname{pqn}(G^{\prime},w^{\prime})=1$ for any given edge-weight function $w^{\prime}$. We construct a PQ-layout $\Gamma$ of $\langle G,w\rangle$ using one priority queue where $w(e)=w^{\prime}(e)$ except for $w(e_{0})=W<\min_{e\in E\setminus\{e_{0}\}}w(e)$. Assuming $u\prec v$ in $\Gamma$, we construct a PQ-layout $\Gamma^{\prime}$ of $\langle G^{\prime},w^{\prime}\rangle$: We place vertex $x_{uv}$ obtained from contracting $e_{0}$ at the position of $u$ in $\Gamma$. Vertices preceding $u$ or succeeding $v$ in $\Gamma$ retain their positions in $\Gamma^{\prime}$, and vertices between $u$ and $v$ in $\Gamma$ are sorted according to the weight of their edges shared with $x_{uv}$. See Appendix B for details.
Corollary 4.3.
The family of graphs with priority queue number 1 is minor-closed.
@@ 4.2 Graphs with Priority Queue Number 1
Proof 4.5.
We describe an algorithm that constructs a PQ-layout $\Gamma$ of $T$; see Figure 3 for an example. For each vertex $v\in V\setminus\{r\}$, we define its weight $w(v)$ as $w(v)=w(vp(v))$. As an initial layout, we first place $r$ as the leftmost vertex and the children of $r$ in increasing order of their weights to the right of $r$. During our algorithm, we keep track of the last expanded vertex $v^{\star}$, which is the rightmost vertex whose children (if any) have already been placed. Note that initially $v^{\star}=r$. We now maintain the following invariants:
@@
Figure 3: A PQ-layout of a weighted rooted tree computed with the algorithm in Lemma 4.4.
@@
Proof 4.11 (Proof Sketch).
We remove the heaviest edge $e^{\star}$ of $C(L)$ and all leaves where the weight of the incident edge is greater than $w(e^{\star})$. We lay out the remaining graph using Lemma 4.6 such that the endpoints of $e^{\star}$ are the first and the last vertex. We reinsert $e^{\star}$ and, in increasing order of the weights of their incident edges, we add the initially removed leaves to the right.
Next, we show that we may attach one caterpillar (Lemma 4.12) or two caterpillars to a single quadrangle (Lemma 4.14) or a single triangle (Corollary 4.16).
@@
Proof 4.13.
We first draw $C$ and $O$ separately using Lemmas 4.6 and 4.8, respectively. Then, we identify the two occurrences of $r$ in $C$ and $O$ with each other, which yields our PQ-layout $\Gamma$ of $G$. This does not result in a forbidden configuration since $r$ is the rightmost vertex of $C$ and the leftmost vertex of $O$ and, hence, there is no edge spanning over $r$.
@@
  • A $4$-cycle $\Box=abcd$.
  • A caterpillar $C_{a}$ with underyling path $P(C_{a})$ starting at vertex $a\in\Box$.
  • A caterpillar $C_{c}$ with underyling path $P(C_{c})$ starting at vertex $c\in\Box$.
@@
Proof 4.15 (Proof Sketch).
We arrange the cycle such that only the heaviest edge is co-active with the other three edges. We use Lemma 4.4 for laying out the one caterpillar with designated leftmost vertex and we use Lemma 4.6 for laying out the other caterpillar with designated rightmost vertex. In the appendix, we argue that we can combine these three subgraphs, potentially adding some leaves later on, while still using only one priority queue.
By Lemma 4.1, if $G$ is a minor of a graph as described in Lemma 4.14, then also $\operatorname{pqn}(G)=1$.
Corollary 4.16.
If $G$ consists of a triangle $\triangle=abc$ and two caterpillars $C_{a}$ and $C_{b}$ with underlying paths starting at vertex $a\in\triangle$ and $b\in\triangle$, respectively, then $\operatorname{pqn}(G)=1$.
Finally, we consider graphs that contain more than one cycle. If we exclude disconnected graphs, there exist exactly two such graphs with priority queue number 1, the complete bipartite graph $K_{2,3}$ and the graph obtained from $K_{4}$ by removing an edge. We will later see that these are the only ones.
@@
In addition, by Lemma 4.1, we obtain the following from Lemma 4.17.
Corollary 4.18.
@@ 4.3 Forbidden Minors for Graphs with Priority Queue Number 1
Each of the graphs $\mathcal{F}_{1},\ldots,\mathcal{F}_{8}$ shown in Figure 4 has priority queue number at least two. The numbers at the edges specify edge-weight functions for which one priority queue does not suffice. In the appendix, we prove the following observation:
@@
We now use Lemma 4.19 to prove the main result of this subsection:
@@
For each graph $\mathcal{F}\in\{\mathcal{F}_{1},\ldots,\mathcal{F}_{8}\}$ (see Figure 4), $\operatorname{pqn}(\mathcal{F})>1$ holds.
Proof 4.21 (Proof of Cases $\mathcal{F}_{1}$$\mathcal{F}_{3}$).
To show $\operatorname{pqn}(\mathcal{F})>1$, it suffices to show $\operatorname{pqn}(\mathcal{F},w)>1$ for some edge weighting $w$. We claim that the edge weightings in Figure 4 require more than one priority queue. We consider each of the eight graphs individually and suppose for a contradiction that there is a PQ-layout $\Gamma$ with $\operatorname{pqn}(\Gamma)=1$. We consider $\mathcal{F}_{1}$$\mathcal{F}_{3}$ here and $\mathcal{F}_{4}$$\mathcal{F}_{8}$ in the appendix.
@@
: Consider the triangle $\triangle abd$. By Lemma 4.19, $a$ or $b$ is the last vertex of $\triangle abd$ (on the spine) since $ab$ is its heaviest edge. In the triangle $\triangle bcd$, $cd$ is the heaviest edge, which implies that $c$ or $d$ is the last vertex of $\triangle bcd$. Hence, neither $b$ nor $d$ can be the (overall) rightmost vertex. Since the heaviest edge of the 4-cycle $abcd$ is again $ab$, $a$ is the rightmost vertex. In the triangle $\triangle acd$, $cd$ is the heaviest edge, which implies that $a$ cannot be the rightmost vertex; a contradiction.
@@
: By Lemma 4.19 and symmetry, we may assume that $a$ and $d$ are the last vertices of $\triangle abc$ and $\triangle cde$, respectively. Moreover, assume by symmetry that $a\prec d$, i.e., $c\prec a\prec d$. However, then edge $ab$ ends below the lighter edge $cd$; a contradiction.
@@
: By Lemma 4.19, $c$ or $d$ is the last vertex of the 5-cycle $abcde$. In the 4-cycle $bcde$, however, $b$ or $e$ is the last vertex; a contradiction.

### v3

@@ 1 Introduction
Stack layouts were introduced by Bernhart and Kainen in 1979 [13] with the motivation to study a “quite natural” edge decomposition technique. A stack layout is a linear layout, that is, the vertices are positioned on a line $\ell$ with the edges being embedded on several half-planes, called pages, delimited by $\ell$. In a stack layout, edges embedded on the same page are forbidden to cross, a rule that contributed to their original designation as book embeddings. An example of stack layout on one page is depicted in Figure˜1(b). The minimum number of pages required to obtain a stack layout of a graph is called the stack number. After their introduction, stack layouts attracted notable interest in the field of theoretical computer science due to their capacity to model certain aspects in several domains, including VLSI design [17, 18], fault-tolerant multiprocessing [18, 40], RNA folding [27], and traffic-light control [32]. This led, in particular, to a series of papers investigating upper bounds on the stack number of planar graphs [16, 28, 31] and culminating at an upper bound of four [42]. Bekos et al. [11] and Yannakakis [43] independently proved that this upper bound is tight.
@@
Based on this latter interpretation of stack layouts, Heath, Leighton and Rosenberg [29, 30] proposed to study queue layouts of graphs where the edges occurring on the same page of the linear layout can be represented by a queue. More precisely, edges ending at $v_{i}$ are first dequeued from a queue, after which edges beginning at $v_{i}$ are enqueued during the left-to-right sweep of $\ell$. Thus, edges on the same page are allowed to cross but forbidden to properly nest. An example of a queue layout on one page is shown in Figure˜1(c). Surprisingly, these quite non-planar representations still found many practical applications, including scheduling [14], VLSI [34], and, a decade after their inception, 3D crossing-free graph drawing [20, 23]. Similarly to stack layouts, one is interested in minimizing the number of pages required to obtain a queue layout, which is known as the queue number. However, in contrast to the stack number, the queue number of planar graphs remained elusive for several decades during which only sublinear (but non-constant) upper bounds have been found [1, 5, 6, 21]. The first significant difference between planar and non-planar graphs was only demonstrated in 2019 when it was shown that bounded degree planar graphs have bounded queue number [9], whereas non-planar bounded degree graphs were already known to not have this property [41]. Recently, Dujmović et al. [22] proved that the queue number of planar graphs is bounded by a constant. Minor improvements have since been made [10, 25].
@@
We introduce a model called priority queue layout, or simply PQ-layout. This model utilizes priority queues for the storage of edges assigned to the same page, where the edge weights correspond to the priorities (i.e., keys) for the data structure. As for stack and queue layouts, in a PQ-layout all vertices of the graph lie on a horizontal line $\ell$, and the edges are assigned to different priority queues according to a left-to-right sweep of $\ell$. When an edge $e$ is dequeued, it must have the minimum priority among the edges stored in the same priority queue as $e$. See Figure˜1(d) for an example, and refer to Section˜2 for a formal definition of our model. Note that, unlike stack and queue layouts, our model allows edges in the same page to cross or properly nest, provided that they have suitable weights.
@@ 2 PQ-Layouts
The order of $u$ and $u^{\prime}$ is irrelevant for F.1. For completeness, the resulting three arrangements of the involved vertices are illustrated in Figures˜2(a), 2(b) and 2(c). Note that the forbidden pseudo-nesting in Figure˜2(b) requires $u=u^{\prime}$. In contrast, the symmetric case where $v=v^{\prime}$ is never forbidden. This is a direct consequence of the asymmetry between the insert and pull operations of priority queues. Thus, unlike for stack and queue layouts, the reverse vertex ordering does not necessarily provide a PQ-layout.
@@
Figure 2: (a)–(c) Arrangements in a PQ-layout that correspond to the Forbidden Configuration F.1. (d) A $k$-inversion, a forbidden arrangement in linear layouts with priority queue number k1k-1.
@@
Some of our proofs utilize configurations where every pair of edges must be assigned to distinct priority queues. Let $\langle G,w\rangle$ be an edge-weighted graph, let $\prec$ be a vertex ordering of $G$, and let $k$ be an integer such that $k\geq 1$. We define a $k$-inversion to be a sequence $(e_{1},\ldots,e_{k})$ of $k$ edges in $G$ with pairwise distinct right endpoints $v_{1},\ldots,v_{k}$, respectively, such that (i) $v_{k}\prec\cdots\prec v_{1}$, where $v_{k}$ and $v_{1}$ are called the left and right end, respectively, (ii) the left endpoint of each of $e_{1},\ldots,e_{k}$ appears to the left of $v_{k}$ in $\prec$, and (iii) $w(e_{1})<\cdots<w(e_{k})$. See Figure˜2(d) for an illustration. Any two edges in a $k$-inversion form a forbidden configuration and thus must be assigned to distinct priority queues. In other words, if, for edge weights $w$ of $G$, every vertex ordering of $\langle G,w\rangle$ contains a $k$-inversion, then $\operatorname{pqn}(G,w)\geq k$ and hence $\operatorname{pqn}(G)\geq k$.
@@ 3 Priority Queue Number of Complete Graphs
Theorem˜3.1 also provides an upper bound for the (non-separated) priority queue number of $K_{n,n}$. Next, we give a corresponding linear lower bound showing that $\operatorname{pqn}(K_{n,n})\in\Theta(n)$.
@@
For $K_{n,n}$, we define an edge-weight function $w$ that has, at each vertex, for every $i\in\{1,\ldots,n\}$, exactly one incident edge of weight ii. [^1] In contrast to the proof of Theorem 3.1, we cannot arbitrarily assign weights to edges incident to each vertex $b\in B$ because this may give twice the same edge weight at a vertex of partition $A$. To this end, we partition the edge set into $n$ perfect matchings $M_{1},\ldots,M_{n}$, and, for each edge $e\in M_{i}$, we set $w(e)=i$.
@@
Note that we cannot directly apply Theorem˜3.1 to $\Gamma^{\prime}$ because we have already fixed an edge-weight function. Instead, we analyze the possible distribution of the edge weights incident to vertices in $\Gamma^{\prime}$ by a representation as a black-and-white grid; [^3] A similar, although not identical, representation is used by Alam et al. [35] to prove the mixed page number of $K_{n,n}$. see the appendix for an illustration and a more extensive description. Our grid has $\frac{n}{2}$ columns that represent the vertices of the latter partition (in order from left to right as they appear in $\Gamma^{\prime}$) and it has $n$ rows that represent the edge weights $\{1,\dots,n\}$ (in order from bottom to top). A grid cell in column $i$ and row $j$ is colored black if the $i$-th vertex is incident to an edge of weight $j$, and it is colored white otherwise. Observe that, in this grid, a strictly monotonically decreasing path of black cells having length $k$ is equivalent to a $k$-inversion in $\Gamma^{\prime}$. We can show that there always exists such a path of length $\left\lceil\frac{3-\sqrt{5}}{4}n\right\rceil$: the candidates for being the first black cell of the path are those who do not have another black cell in their top left region of the grid. We find these cells, color them white and iteratively repeat this process to find the candidates for being the second, third, etc. black cell of the path. We can prove that we need at least $\left\lceil\frac{3-\sqrt{5}}{4}n\right\rceil$ iterations until the grid is white.
Since $K_{n}$ contains $K_{\lfloor\frac{n}{2}\rfloor,\lfloor\frac{n}{2}\rfloor}$ as a subgraph, Theorem˜3.3 immediately implies a linear lower bound on $\operatorname{pqn}(K_{n})$:
@@ 4 Characterizing Graphs with Priority Queue Number 1
We study which graphs have priority queue number 1. We will provide a characterization of the graphs with priority queue number 1 in terms of forbidden minors (Section˜4.3). We first describe families of graphs with priority queue number 1 (Section˜4.1). Then, we give a set of eight forbidden minors (Section˜4.2). The characterization of Section˜4.3 follows by proving that every graph that does not contain any of the forbidden minors of Section˜4.2 falls in one of the families of Section˜4.1. This characterization leads to a linear-time recognition algorithm.
@@ 4.1 Graphs with Priority Queue Number 1
Proof 4.2.
We describe an algorithm that constructs a PQ-layout $\Gamma$ of $T$; see Figure˜3 for an example. For each vertex $v\in V\setminus\{r\}$, we define its weight $w(v)$ as $w(v)=w(vp(v))$. As an initial layout, we first place $r$ as the leftmost vertex and the children of $r$ in increasing order of their weights to the right of $r$. During our algorithm, we keep track of the last expanded vertex $v^{\star}$, which is the rightmost vertex whose children (if any) have already been placed. Note that initially $v^{\star}=r$. We now maintain the following invariants:
@@
Figure 3: A PQ-layout of a weighted rooted tree computed with the algorithm in Lemma˜4.1.
@@
Proof 4.8 (Proof Sketch).
We remove the heaviest edge $e^{\star}$ of $C(L)$ and all leaves where the weight of the incident edge is greater than $w(e^{\star})$. We lay out the remaining graph using Lemma˜4.3 such that the endpoints of $e^{\star}$ are the first and the last vertex. We reinsert $e^{\star}$ and, in increasing order of the weights of their incident edges, we add the initially removed leaves to the right.
Next, we show that we may attach one caterpillar to any cycle (Lemma˜4.9) or two caterpillars to a single triangle or quadrangle (Lemma˜4.11).
@@
Proof 4.10.
We first draw $C$ and $O$ separately using Lemmas˜4.3 and 4.5, respectively. Then, we identify the two occurrences of $r$ in $C$ and $O$ with each other, which yields our PQ-layout $\Gamma$ of $G$. This does not result in a forbidden configuration since $r$ is the rightmost vertex of $C$ and the leftmost vertex of $O$ and, hence, there is no edge spanning over $r$.
@@
  • A 3-cycle $\triangle=abc$ or a 4-cycle $\Box=abcd$.
  • A caterpillar $C_{a}$ with underyling path $P(C_{a})$ starting at vertex $a$.
  • A caterpillar $C_{c}$ with underyling path $P(C_{c})$ starting at vertex $c$.
@@
Proof 4.12 (Proof Sketch).
We arrange the cycle such that only the heaviest edge is co-active with the other edges of the cycle. We use Lemma˜4.1 for laying out the one caterpillar with designated leftmost vertex and we use Lemma˜4.3 for laying out the other caterpillar with designated rightmost vertex. In the appendix, we argue that we can combine these three subgraphs, potentially adding some leaves later on, while still using only one priority queue.
Finally, we consider graphs that contain more than one cycle. If we exclude disconnected graphs, there exist exactly two such graphs with priority queue number 1, the complete bipartite graph $K_{2,3}$ and the graph obtained from $K_{4}$ by removing an edge. We will later see that these are the only ones. The proofs that one priority queue suffices can be found in the appendix.
@@
[…]
@@ 4.2 Forbidden Minors for Graphs with Priority Queue Number 1
Next, we prove that each of the graphs $\mathcal{F}_{1},\ldots,\mathcal{F}_{8}$ shown in Figure˜4 has priority queue number at least two. The numbers at the edges specify edge-weight functions for which one priority queue does not suffice. Moreover, we prove that every graph that has one of $\mathcal{F}_{1},\ldots,\mathcal{F}_{8}$ as a minor, has priority queue number at least two, too. In the appendix, we prove the following helpful observation:
@@
We now use Lemma˜4.15 to prove the main result of this subsection:
@@
For every graph $G$ that has some of $\mathcal{F}_{1},\ldots,\mathcal{F}_{8}$ (see Figure˜4) as a minor, $\operatorname{pqn}(G)>1$ holds.
Proof 4.17 (Proof of Cases $\mathcal{F}_{1}$$\mathcal{F}_{3}$).
To show $\operatorname{pqn}(G)>1$, it suffices to show $\operatorname{pqn}(G,w)>1$ for some edge weighting $w$. For each $\mathcal{F}\in\{\mathcal{F}_{1},\ldots,\mathcal{F}_{8}\}$, we prove that $\mathcal{F}$ itself has $\operatorname{pqn}(\mathcal{F})>1$ and we prove that for any reversal of an edge contractions the resulting graph $G$ has $\operatorname{pqn}(G)>1$. In a reversal of an edge contraction, we split a vertex $v$ into two vertices $v_{1}$ and $v_{2}$, add the edge $v_{1}v_{2}$, and assign the edges incident to $v$ arbitrarily to $v_{1}$ or $v_{2}$ or both. We call the reversal of an edge contraction a vertex split. Note that it suffices to consider vertex splits since taking a supergraph of $\mathcal{F}$ can never reduce the number of priority queues needed for the edges of $\mathcal{F}$ alone. Further note that is suffices to consider vertex splits where each edge incident to $v$ is assigned to exactly one of $v_{1}$ and $v_{2}$ and each of $v_{1}$ and $v_{2}$ gets at least one incident edge of $v$. If in $G$ this was not the case, we could remove duplicates of incident edges of $v$ and we could remove the one of $v_{1}$ and $v_{2}$ that has degree 1; the resulting subgraph $G^{\prime}$ of $G$ would have $\mathcal{F}$ as a minor.
We claim that the edge weightings in Figure˜4 require more than one priority queue. We consider each of the eight graphs and the graphs for which they are minors individually and suppose for a contradiction that there is a PQ-layout $\Gamma$ of $G$ with $\operatorname{pqn}(\Gamma)=1$. We consider $\mathcal{F}_{1}$$\mathcal{F}_{3}$ here and $\mathcal{F}_{4}$$\mathcal{F}_{8}$ in the appendix.
@@
: First consider $G=\mathcal{F}_{1}$. By Lemma˜4.15, $c$ or $d$ is the last vertex of the 5-cycle abcdeabcde. In the 4-cycle bcdebcde, however, $b$ or $e$ is the last vertex; a contradiction.
Now let $G$ be any graph that has $\mathcal{F}_{1}$ as a minor. As explained above, we do not need to consider vertex splits where new leaves or cycles arise. So, with a sequence of vertex splits, we can increase (i) the length of the path bcdebcde, (ii) the length of the path baebae, or (iii) the length of the path $be$. In any case, we assign any new edge a weight of 1. Then, our arguments about the last vertices also apply for larger cycles.
@@
: First consider $G=\mathcal{F}_{2}$. By Lemma˜4.15 and symmetry, we may assume that $a$ and $d$ are the last vertices of $\triangle abc$ and $\triangle cde$, respectively. Moreover, assume by symmetry that $a\prec d$, i.e., $c\prec a\prec d$. However, then edge $ab$ ends below the lighter edge $cd$; a contradiction.
Now let $G$ be any graph that has $\mathcal{F}_{2}$ as a minor. With a sequence of vertex splits, we can (i) increase the length of the cycles abcabc or cdecde, (ii) create a path (replacing $c$) between the two cycles; we denote the two endpoints of that path as $c_{1}$ and $c_{2}$ where $c_{1}$ is part of the one cycle and $c_{2}$ is part of the other cycle. In any case, we assign any new edge a weight of 1. In (i), we can still assume that $c\prec a\prec d$. In contrast to before $c$ might not be a neighbor of $a$ or $d$ any more. However, there is a path of edges with weight 1 between $c$ and $a$ and between $c$ and $d$. Some edge $e$ on the path between $c$ and $d$ spans over $a$, which is the endpoint of the heavier edge $ab$; a contradiction. In (ii), we establish again (by Lemma˜4.15 and symmetry up to the length of the cycles, which is not relevant here) that $c_{1}\prec a$, $c_{2}\prec d$, $a\prec d$. To obtain the contradiction that an edge of weight 1 on the path between $c_{2}$ and $d$ spans over the endpoint of the edge $ab$ of weight 2, we show that $c_{2}\prec a$. Suppose for a contradiction that $c_{1}\prec a\prec c_{2}$. Then, there is some edge of weight 1 on the path between $c_{1}$ and $c_{2}$ that spans over the heavier edge $ab$; a contradiction.
@@
: First consider $G=\mathcal{F}_{3}$. Consider the triangle $\triangle abd$. By Lemma˜4.15, $a$ or $b$ is the last vertex of $\triangle abd$ (on the spine) since $ab$ is its heaviest edge. In the triangle $\triangle bcd$, $cd$ is the heaviest edge, which implies that $c$ or $d$ is the last vertex of $\triangle bcd$. Hence, neither $b$ nor $d$ can be the (overall) rightmost vertex. Since the heaviest edge of the 4-cycle abcdabcd is again $ab$, $a$ is the rightmost vertex. In the triangle $\triangle acd$, $cd$ is the heaviest edge, which implies that $a$ cannot be the rightmost vertex; a contradiction.
Now let $G$ be any graph that has $\mathcal{F}_{3}$ as a minor. Since $\mathcal{F}_{3}$ is the complete graph on four vertices (i.e., $K_{4}$), it is highly symmetric and the first vertex split makes one of the four 3-cycles a 4-cycle. This new graph is precisely $\mathcal{F}_{1}$ with one additional edge.